# Evolutionary Retrosynthetic Route Planning

### v1

@@ Abstract
Molecular retrosynthesis is a significant and complex problem in the field of chemistry, however, traditional manual synthesis methods not only need well-trained experts but also are time-consuming. With the development of big data and machine learning, artificial intelligence (AI) based retrosynthesis is attracting more attention and is becoming a valuable tool for molecular retrosynthesis. At present, Monte Carlo tree search is a mainstream search framework employed to address this problem. Nevertheless, its search efficiency is compromised by its large search space. Therefore, we propose a novel approach for retrosynthetic route planning based on evolutionary optimization, marking the first use of Evolutionary Algorithm (EA) in the field of multi-step retrosynthesis. The proposed method involves modeling the retrosynthetic problem into an optimization problem, defining the search space and operators. Additionally, to improve the search efficiency, a parallel strategy is implemented. The new approach is applied to four case products, and is compared with Monte Carlo tree search. The experimental results show that, in comparison to the Monte Carlo tree search algorithm, EA significantly reduces the number of calling single-step model by an average of 53.9%. The time required to search three solutions decreased by an average of 83.9%, and the number of feasible search routes increases by 5 times. The source code is available at https://github.com/ilog-ecnu/EvoRRP.
@@ I Introduction
Molecular retrosynthesis is of great importance in various fields[1, 2], including drug synthesis and catalyst design. It enables rapid determination of potential pathways and starting materials for the synthesis of target compounds, while also providing essential guidance for synthesis routes. Retrosynthetic process is divided into two processes: single-step retrosynthesis and multi-step search process. The single-step retrosynthesis process involves breaking down an organic molecule into its original reactants. As illustrated in Fig. 1, the retrosynthetic route for target product is generated using this method. The intermediate molecules, $R^{1}$ and $W^{1}$, are obtained through the single-step model with the input target product $P$. $W^{1}$ belongs to the set $\Psi$, which represents the building block dataset containing commercially available molecules, serving as the terminal reactant database. The process continues by selecting $R^{1}$ as input for the single-step model, and this cycle repeats until all products belong to $\Psi$, forming a multi-step retrosynthetic reactions.
@@
In recent years, neural network (NN), particularly architectures like transformers, has played a crucial role in advancing this field. A transformer architecture[3] has become a common choice for single-step retrosynthesis tasks. It has achieved significant advancements and greatly improved search effectiveness, leading to remarkable progress in the field of retrosynthesis. Currently, researchers are focusing on multi-step retrosynthetic route planning[4]. This involves a combination of single-step retrosynthetic models based on transformer or convolutional neural network, along with a multi-step Monte Carlo tree search framework. This approach has emerged as a valuable asset in the field of chemistry. It empowers chemists to expedite the synthesis of novel compounds with enhanced efficiency and precision, opening new vistas of exploration and discovery in the vast realm of chemical synthesis.
@@
Monte Carlo Tree Search (MCTS) is a decision-making algorithm initially applied to multi-step retrosynthesis by Marwin et al. [1]. They used a neural network for single-step retrosynthesis, which is an end-to-end sequence generation effort, and MCTS for route search effort. The neural network was trained on a large dataset of known organic reactions to predict the outcome of a given chemical reaction. MCTS was used to search for the optimal synthetic route by exploring the possible chemical transformations that could be applied to the target molecule, evaluating the likelihood of each transformation using the neural network, and selecting the best one to apply. The advantage of this approach is that it can efficiently explore a large search space and find a good solution, while the neural network can provide accurate predictions of chemical reactions. Lin et al.[3] also applied MCTS with a heuristic scoring function for multi-step retrosynthetic route planning. MCTS is a powerful algorithm for decision making in complex problems, however, its effectiveness and efficiency depend on several factors such as the complexity of the problem, the branching factor of the tree, the number of simulations required, and the computational resources available. Therefore, researchers must carefully design the algorithm and its parameters to achieve the desired results. While MCTS has many advantages in decision-making problems with a large number of possible actions and states, there are also a few drawbacks to consider. Firstly, it may not work well for games or problems that have specific structures or constraints. Secondly, the search efficiency of MCTS is not very high due to its requirement of a large search space during the exploration process. Lastly, MCTS can be computationally expensive, particularly when the branching factor of the tree is high or when a large number of simulations are needed. While MCTS can be parallelized[5], it is not commonly used in the context of retrosynthetic route planning.
@@
In summary, we propose a method for retrosynthetic route planning using evolutionary optimization, which can purposefully find the optimal solution while adhering to the constraints of the objective function. The method defines the search space and limits the search scope, thereby reducing the generation of infeasible solutions. Moreover, the proposed method employs parallel computation to reduce the search time and improve search efficiency and the number of search routes. The contributions of this paper can be summarized as follows:
  • We introduce a novel method for retrosynthetic problem, which can be solved by Evolutionary Algorithm (EA), marking the first use of EA in the field of multi-step retrosynthesis.
  • The proposed method greatly reduces the single-step model calls, decreasing the frequency of generating invalid solutions, and improving search efficiency.
  • Additionally, a parallel strategy is implemented to improve the search efficiency during the searching process.
@@ II-A Single-step retrosynthesis
Single-step retrosynthesis can be categorized into template-based[6, 7, 8], template-free[9, 10, 3, 11], and semi-template[12, 13, 14] methods. Template-based retrosynthesis[15] employs a pre-defined set of reaction templates to guide retrosynthetic analysis. These templates are based on frequently used organic synthesis reactions and can suggest possible starting materials for a desired target molecule. For instance, if the target molecule has a carbonyl group, the template-based approach may recommend using a nucleophilic addition reaction to introduce the carbonyl group and then working backwards to identify the precursor molecules for the reaction. Though this approach yields high accuracy, it requires a considerable amount of computation. Furthermore, the rule cannot cover the response range completely, and its scalability is limited. Template-based retrosynthetic reactions utilize reaction rules that may lead to reactivity conflicts. To deal with this, deep neural networks[16] have been used to resolve this problem. Computer-assisted synthesis planning (CASP)[17] was also gaining attention and molecular similarity has been found to be an effective metric for proposing and ranking single-step retrosynthetic disconnections. Deep learning-based method[18], which combine local reactivity and global attention mechanisms, was commonly used for retrosynthetic reaction prediction.
@@
Template-free retrosynthesis[19] is a more flexible approach that does not rely on pre-defined reaction templates. Instead, it involves breaking down the target molecule into smaller fragments based on the functional groups present and then considering possible ways to connect these fragments. This open-ended approach can be useful for exploring a wider range of possible synthetic pathways. However, the accuracy of this approach is not high, especially when the reaction type is unknown, although its scalability is good. Template-free methods, on the other hand, have high scalability but lower accuracy compared to template-based methods. Retroformer[20] used a novel differentiable MCTS method to directly predict products from reactants, achieving end-to-end retrosynthesis. RetroPrime[21] combined global and local contextual information, employed a multi-layer transformer architecture, and used a multi-task learning strategy to improve the accuracy and robustness of retrosynthesis prediction. Besides transformer, graph-based truncated attention (GTA) model[22] employed a truncated attention mechanism to handle interactions between different nodes and combined a reaction library with graph neural networks for retrosynthetic prediction.
@@
Several semi-template-based approaches[23, 24] also exist, such as a molecular graph-enhanced transformer model[25] that represented molecular structures in graph form and used graph-based self-attention mechanisms to handle interactions between different atoms in the molecule. Additionally, the model used an adaptive structural embedding method to enhance its ability to model molecular structures.
@@ II-B Multi-step search process
Multi-step retrosynthetic reaction involves a series of consecutive single-step retrosynthesis to gradually break down the target molecule, starting from the target molecule and undergoing transformations of multiple intermediate compounds, ultimately yielding simpler starting materials. It emphasizes the requirement for multiple reaction steps to accomplish the decomposition of the target molecule, with each step serving as a crucial component of the retrosynthesis. One such approach is the game tree search[26], which generated new drug molecules by searching for potential reaction paths. This method combined traditional rule-based and deep learning methods and used a reinforcement learning algorithm, called game tree search, to guide the drug molecule construction process. Another method is the one proposed by Klucznik et al.[27], which planned the synthesis path and predicted possible side reactions based on the structural characteristics and synthetic difficulty of the target compounds. However, this approach had poor performance in terms of speed of synthesis. To solve this issue, a computer program[2] was developed to predict the optimal reaction conditions for target compounds, which was then used to control the robotic platform to perform multiple reactions simultaneously in flow reactors. Retro*[28] combined neural networks with MCTS to generate synthetic pathways. Additionally, a computational framework[29], integrating a reaction database, a generative model, and an evaluation function, was developed to guide the MCTS-based search for green synthetic pathways. The Reinforcement Learning (RL) algorithm was employed to learn from previous MCTS searches and update the evaluation function to improve search efficiency. These methods demonstrate the potential of computer-assisted synthesis planning in improving the efficiency and accuracy of chemical reactions.
@@ II-C Evolutionary optimization
Evolutionary algorithm (EA) is a kind of population-based optimization algorithm, through the selection of the fittest among individuals to select offspring, so as to solve some complex optimization problems. There are several operators that can be used to generate new solutions, such as genetic algorithm (GA)[30], the differential evolution (DE)[31], the particle swarm optimization (PSO)[32]. Other approaches include evolution strategies (ES)[33], evolutionary programming (EP)[34, 35] and genetic programming (GP)[36], and estimation of distribution algorithms (EDA)[37]. EDA used machine learning methods to generate new solutions and can converge to the global optimum faster than traditional EA frameworks. EDA can be classified into three types: univariate EDA[38], multivariate EDA[39, 40, 41, 42, 43, 44] and multi-objective EDA[45, 46, 47, 48]. Population-Based Incremental Learning (PBIL)[49] was a representative univariate EDA that combined genetic search-based function optimization and competitive learning. It used a probabilistic model to generate candidate solutions and encouraged diversity by introducing random perturbations. Pelikan et al.[50] proposed a probabilistic modeling approach, a kind of multivariate EDA method, based on Bayesian networks to solve the linkage problem and estimate the joint probability distribution of variables in a high-dimensional search space. BMDA[51] was another well-known multivariate EDA method that estimates the joint probability distribution of decision variables and used this information to guide the search for the global optimum. Naive MIDEA[52] was a multi-objective optimization algorithm that balanced convergence and diversity by using a fitness assignment scheme based on the distance between solutions. It also incorporated a crowding distance measure to promote diversity in the population. A hybrid Bayesian optimization algorithm[53] was proposed to combine Gaussian process with tree-structure Bayesian optimization and introduce the variance adaptive mechanism, which can solve high-dimensional complex function optimization problems. In summary, EDA has the ability to learn the problem structure and is suitable for solving real optimization problems with special situations. There are also some typical application cases[54, 55, 56] of evolutionary optimization. For instance, In the study by Weber et al.[56], they proposed a method to process multi-component reactions using evolutionary search methods for drug discovery. Multi-component reactions (MCRs) offer a novel method for synthesizing a wide range of compounds and compound libraries efficiently. Once considered merely a chemistry curiosity, MCRs are now acknowledged for their growing significance in drug discovery, particularly in lead discovery and optimization. This evolutionary search method primarily emphasizes in silico filtering, but one of its limitations is the lack of integration with deep learning techniques.

### v2

@@ Abstract
Molecular retrosynthesis is a significant and complex problem in the field of chemistry, however, traditional manual synthesis methods not only need well-trained experts but also are time-consuming. With the development of big data and machine learning, artificial intelligence (AI) based retrosynthesis is attracting more attention and has become a valuable tool for molecular retrosynthesis. At present, Monte Carlo tree search is a mainstream search framework employed to address this problem. Nevertheless, its search efficiency is compromised by its large search space. Therefore, this paper proposes a novel approach for retrosynthetic route planning based on evolutionary optimization, marking the first use of Evolutionary Algorithm (EA) in the field of multi-step retrosynthesis. The proposed method involves modeling the retrosynthetic problem into an optimization problem, defining the search space and operators. Additionally, to improve the search efficiency, a parallel strategy is implemented. The new approach is applied to four case products and compared with Monte Carlo tree search. The experimental results show that, in comparison to the Monte Carlo tree search algorithm, EA significantly reduces the number of calling single-step model by an average of 53.9%. The time required to search three solutions decreases by an average of 83.9%, and the number of feasible search routes increases by 1.38 times. The source code is available at https://github.com/ilog-ecnu/EvoRRP.
@@ I Introduction
Molecular retrosynthesis is of great importance in various fields[1, 2], including drug synthesis and catalyst design. It enables rapid determination of potential pathways and starting materials for the synthesis of target compounds, effectively providing essential guidance for synthesis routes. Retrosynthesis contains two types of process: single-step retrosynthesis and multi-step search process. The single-step retrosynthesis process involves breaking down an organic molecule into its original reactants. As illustrated in Fig. 1, the retrosynthetic route for target product is generated using this single-step method. Moreover, the intermediate molecules, $R^{1}$ and $W^{1}$, are obtained through the single-step model with the input target product $P$. $W^{1}$ belongs to the set $\Psi$, which represents the building block dataset containing commercially available molecules, serving as the terminal reactant database. The process continues by selecting $R^{1}$ as input for the single-step model, and this cycle repeats until all products belong to $\Psi$, forming a multi-step retrosynthetic reactions.
@@
In recent years, single-step model, particularly used in architectures like transformers, has played a crucial role in advancing this field. A transformer architecture[3] has become a common choice and practical solution for single-step retrosynthesis tasks. It has achieved significant advancements and greatly improved search effectiveness, leading to remarkable progress in the field of retrosynthesis. Currently, researchers are focusing on multi-step retrosynthetic route planning[4]. This involves a combination of single-step retrosynthetic models based on transformer or convolutional neural network, along with a multi-step Monte Carlo tree search (MCTS) framework. This approach has emerged as a valuable asset in the field of chemistry. It empowers chemists to expedite the synthesis of novel compounds with enhanced efficiency and precision, opening new vistas of exploration and discovery in the vast realm of chemical synthesis.
@@
Among multi-step search frameworks, MCTS is a decision-making algorithm initially applied to multi-step retrosynthesis by Marwin et al. [1]. They used a neural network for single-step retrosynthesis, which is an end-to-end sequence generation effort, and MCTS for route search effort. The neural network was trained on a large dataset of known organic reactions to predict the outcome of a given chemical reaction. MCTS was used to search for the optimal synthetic route by exploring the possible chemical transformations that could be applied to the target molecule, evaluating the likelihood of each transformation using the neural network and selecting the best one to apply. The advantage of this approach is that it can efficiently explore a large search space and find a good solution, while the neural network can provide accurate predictions of chemical reactions. Lin et al.[3] also applied MCTS with a heuristic scoring function for multi-step retrosynthetic route planning. MCTS is a powerful algorithm for decision making in complex problems, however, its effectiveness and efficiency depend on several factors, such as the complexity of the problem, the branching factor of the tree, the number of simulations required, and the computational resources available. Therefore, researchers must carefully design the algorithm and its parameters to achieve the desired results. While MCTS has many advantages in decision-making problems with a large number of possible actions and states, there are also a few drawbacks to consider. Firstly, it may not work well for games or problems that have specific structures or constraints. Secondly, the search efficiency of MCTS is not very high due to its requirement of a large search space during the exploration process. Lastly, MCTS can be computationally expensive, particularly when the branching factor of the tree is high or when a large number of simulations are needed. MCTS can be parallelized[5], however, it is not commonly used in the context of retrosynthetic route planning.
@@
Although MCTS has performed well on multi-step retrosynthesis from a decision-making point of view, its search efficiency is compromised by its large search space. Therefore, this paper proposes a method for retrosynthetic route planning using evolutionary optimization, which can purposefully find solutions while adhering to the constraints of the objective function. Secondly, the method defines the search space and limits the search scope, thereby reducing the generation of infeasible solutions. Thirdly, the proposed method employs parallel computation to reduce the search time and improve search efficiency and the number of search routes. The contributions of this paper can be summarized as follows:
  • Evolutionary Algorithm (EA) is used for solving retrosynthetic problem, marking the first use of EA in the field of multi-step retrosynthesis.
  • The single-step model is utilized to expand the tree nodes in the process of retrosynthetic route planning. The proposed method greatly reduces the single-step model calls, decreasing the frequency of generating invalid solutions, and improving search efficiency.
  • Since each individual in the population is independent of each other, a parallel-EA is implemented to improve the search efficiency during the searching process.
@@ II Related work
This section describes three aspects, namely, single-step retrosynthesis, multi-step search process, and evolutionary optimization. The single-step retrosynthesis is analyzed from three categories. The multi-step search process describes some state-of-the-art methods. Evolutionary optimization introduces some common genetic operators, classical probabilistic models, and some practical applications.
@@ II-A Single-step retrosynthesis
Single-step retrosynthesis can be categorized into template-based, template-free, and semi-template methods. Template-based retrosynthesis[6] employs a pre-defined set of reaction templates to guide retrosynthetic analysis. These templates are based on frequently used organic synthesis reactions and can suggest possible starting materials for a desired target molecule. For instance, if the target molecule has a carbonyl group, the template-based approach may recommend using a nucleophilic addition reaction to introduce the carbonyl group and then working backwards to identify the precursor molecules for the reaction. Although this approach yields high accuracy, it requires a considerable amount of computation. Furthermore, the rule cannot cover the response range completely, and its scalability is limited. Template-based retrosynthetic reactions utilize reaction rules that may lead to reactivity conflicts. To deal with this, deep neural networks[7] have been used to resolve this problem. Computer-assisted synthesis planning (CASP)[8] was also gaining attention and molecular similarity has been found to be an effective metric for proposing and ranking single-step retrosynthetic disconnections. Deep learning-based method[9], which combine local reactivity and global attention mechanisms, was commonly used for retrosynthetic reaction prediction.
@@
Template-free retrosynthesis[10] is a more flexible approach that does not rely on pre-defined reaction templates. Instead, it involves breaking down the target molecule into smaller fragments based on the functional groups present and then considering possible ways to connect these fragments. This open-ended approach can be useful for exploring a wider range of possible synthetic pathways. However, the accuracy of this approach is not high, especially when the reaction type is unknown, although its scalability is good. Template-free methods, on the other hand, have high scalability but lower accuracy compared to template-based methods. Retroformer[11] used a novel differentiable MCTS method to directly predict products from reactants, achieving end-to-end retrosynthesis. RetroPrime[12] combined global and local contextual information, employed a multi-layer transformer architecture, and used a multi-task learning strategy to improve accuracy and robustness of retrosynthesis prediction. Beside transformer, graph-based truncated attention (GTA) model[13] employed a truncated attention mechanism to handle interactions between different nodes and combined a reaction library with graph neural networks for retrosynthetic prediction.
@@
Several semi-template-based approaches also exist, such as a molecular graph-enhanced transformer model[14] that represented molecular structures in graph form and used graph-based self-attention mechanisms to handle interactions between different atoms in the molecule. Additionally, the model used an adaptive structural embedding method to enhance its ability to model molecular structures.
@@ II-B Multi-step search process
Multi-step retrosynthetic reaction involves a series of consecutive single-step retrosynthesis to gradually break down the target molecule, starting from the target molecule and undergoing transformations of multiple intermediate compounds, ultimately yielding simpler starting materials. It emphasizes the requirement for multiple reaction steps to accomplish the decomposition of the target molecule, with each step serving as a crucial component of the retrosynthesis.
The state-of-the-art multi-step retrosynthetic methods include beam search, A* algorithm, MCTS, and so on. Schwaller et al.[15] introduced a single-step retrosynthetic model predicting reactants, and the optimal synthetic pathway was found through a beam search. Meanwhile, Retro*[16], based on A* algorithm, combined neural networks to generate synthetic pathways efficiently. In addition to these two methods, there is another method based on MCTS search, a computational framework[17], integrating a reaction database, a generative model, and an evaluation function, was developed to guide the MCTS-based search for green synthetic pathways. The Reinforcement Learning (RL) algorithm was employed to learn from previous MCTS searches and update the evaluation function to improve search efficiency. Game tree search[18] generated new drug molecules by searching for potential reaction paths. This method combined traditional rule-based with deep learning methods and used a reinforcement learning algorithm to guide the drug molecule construction process. Another method is the one proposed by Klucznik et al.[19], which planned the synthesis path and predicted possible side reactions based on the structural characteristics and synthetic difficulty of the target compounds. However, this approach had poor performance in terms of speed of synthesis. To deal with this issue, a computer program[2] was developed to predict the optimal reaction conditions for target compounds, which was then used to control the robotic platform to perform multiple reactions simultaneously in flow reactors.
Among the above methods, beam search and A* algorithm are relatively minority algorithms. While deep learning and reinforcement learning are more effective in solving retrosynthetic problems, they are not particularly suitable due to the nature of the algorithm. MCTS[1] stands out on retrosynthetic problems, especially for such single-step model and multi-step search combinations. Therefore, MCTS is a more general and representative retrosynthetic route search algorithm. These methods demonstrate the potential of computer-assisted synthesis planning in improving the efficiency and accuracy of chemical reactions.
@@ II-C Evolutionary optimization
Evolutionary algorithm (EA) is a kind of population-based optimization algorithm, through the selection of the fittest among individuals to select offspring, so as to deal with some complex optimization problems. Evolutionary algorithms can be broadly divided into three categories: genetic based, probabilistic model-based, and individual coding based approaches. The first category includes genetic algorithm (GA)[20], differential evolution (DE)[21], and particle swarm optimization (PSO)[22]. The second one is a kind of probabilistic model-based approaches. Yao et al.[23] used Binary Genetic Algorithm (BGA) to the randomized algorithm to generate the initial population. During the optimization process of the algorithm, Bayesian network was used to do a local search and a depth-first algorithm was used to break the loop. Lee et al.[24] presented a new approach for learning the structure of Bayesian networks using a dual genetic algorithm. This method considered both the topology and ordering of BN nodes, expanding the solution space. Corriveau et al.[25] proposed a Bayesian network-based adaptive framework, BNGA, for addressing parameter setting dependencies in a steady-state genetic algorithm. The third type is based on individual coding, such as binary coding, real coding, string coding, tree coding, graph coding and so on. Other approaches include evolution strategies (ES)[26], evolutionary programming (EP)[27, 28] and genetic programming (GP)[29], and estimation of distribution algorithms (EDA)[30].
@@
EDA uses machine learning methods to generate new solutions and can converge to the global optimum faster than traditional EA frameworks. It can be classified into three types: univariate EDA, multi-variate EDA, and multi-objective EDA. Population-Based Incremental Learning (PBIL)[31] is a representative univariate EDA that combines genetic search-based function optimization and competitive learning. It uses a probabilistic model to generate candidate solutions and encourages diversity by introducing random perturbations. EDA/LS[32], a kind of multi-variate EDA method, samples new solutions by constructing a probabilistic model as a genetic operator. This method can assign a higher generation probability in the dominant region and accelerate the algorithm convergence. CMA-ES[33] is also a very classical multi-variate method using probabilistic model for sampling, and also a common probabilistic gradientless optimization method. The proposed method chooses EDA/LS (referred to EDA in the following context) as the search framework. Pelikan et al.[34] proposed a probabilistic modeling approach, a kind of multivariate EDA method, based on Bayesian networks to solve the linkage problem and estimate the joint probability distribution of variables in a high-dimensional search space. BMDA[35] is another well-known multivariate EDA method that estimates the joint probability distribution of decision variables and uses this information to guide the search for the global optimum. Naive MIDEA[36] is a multi-objective optimization algorithm that balances convergence and diversity by using a fitness assignment scheme based on the distance between solutions. It also incorporates a crowding distance measure to promote diversity in the population. Recently, some regularity modeling approaches, i.e., RM-MEDA [37] and its variants [38], have been proposed to approximate the manifold structure of Pareto optimal solutions of multiobjective optimization problems. In summary, EDA has the ability to learn the problem structure and is suitable for solving real optimization problems with special situations.
There are also some typical application cases[39, 40] of evolutionary optimization. For instance, In the study by Weber et al.[41], they proposed a method to process multi-component reactions using evolutionary search methods for drug discovery. Multi-component reactions (MCRs) offer a novel method for synthesizing a wide range of compounds and compound libraries efficiently. Once considered merely a chemistry curiosity, MCRs are now acknowledged for their growing significance in drug discovery, particularly in lead discovery and optimization. This evolutionary search method primarily emphasizes in silico filtering, but one of its limitations is the lack of integration with deep learning techniques. Another practical application, called MemPBPF, was developed by Orozco-Rosas et al.[42] using a combination of membrane computing, a pseudo-bacterial genetic algorithm, and the artificial potential field method. The genetic algorithm component was utilized to evolve the necessary parameters within the artificial potential field method. This approach enabled the algorithm to iteratively optimize the parameters and find the best settings for generating feasible and safe paths for autonomous mobile robots. Also proposed by Orozco-Rosas et al.[43] is the memEAPF approach, which combines membrane computing, a genetic algorithm, and the artificial potential field method for mobile robot path planning. The approach evolves parameters within delimited compartments to minimize path length. Parallel implementations demonstrate its efficiency.
After the realization of retrosynthetic route planning algorithm, researchers can continue to study from the practical application point of view, such as Visual product[44], an industrial robotic manipulator[45] and a practical issue like machine modelling[46] can be considered.